OpenStudy (anonymous):

Please help I don't understand this at all: Approximate the real number solution(s) to the polynomial function f(x) = x3 + 4x2 + x − 6?

OpenStudy (anonymous):

Will become fan and add medal

Nnesha (nnesha):

lol

OpenStudy (anonymous):

Hi @Nnesha, do you think you can help?

OpenStudy (pawanyadav):

We don't have a perticular formula for checking solution of a cubic polynomial. So we need to check it... Start with smallest factor of 6 ...X=1 ,if you get f(1) =0 ,than 1 is its solution

Nnesha (nnesha):

looks like p/q method would work

OpenStudy (anonymous):

i think i have to find the factors of -6 and put them over the factors of 1 right?

Nnesha (nnesha):

\[\rm \frac{ p }{ q }=\frac{ \pm factors~ of~constant }{ \pm factors~of ~leading~coefficient }\]

OpenStudy (pawanyadav):

Yes...

Nnesha (nnesha):

yep that's correct and then test each possible solution by either using synthetic division or replacing x for its value

OpenStudy (anonymous):

so i would end up with \[\frac{ 3 }{ 1 }, \frac{ 2 }{ 1 }, \frac{ 6 }{ 1 }, \frac{ 1 }{ 1 }\]

Nnesha (nnesha):

it should be plus/minus

OpenStudy (anonymous):

of course sorry i forgot

OpenStudy (anonymous):

then i would plug either +/-3, +/-2, +/- 6, +/-1 into my function and observe which number results in f(x) = 0. Right?

Nnesha (nnesha):

yep and remember if there are duplicates numbers after dividing p/q u should eliminate it

OpenStudy (pawanyadav):

Here, f(1)=0 Means (x-1) is its factor. You can find remaining factor by dividing f(x) by (x-1). You will get a quadratic equation And you know how to solve them.

OpenStudy (anonymous):

so would i synthetically divide x-1 by x^3 + 4x^2 + x - 6 (I'm terrible at math so i apologize)

OpenStudy (pawanyadav):

Yes,

Nnesha (nnesha):

or you can find \[\rm f(\pm1),f(\pm2) f(\pm 3),f(\pm6)\] and see which one give you 0 as final answer that would be the real zero ~zeros (x-intercept) is a point where the graph intersect the x-axis when y=0 and in this question f(x) is same as y

OpenStudy (anonymous):

Ill try both your methods although the second seems more time consuming so ill try synthetic division first

Nnesha (nnesha):

true that if you are familiar with the synthetic division method then ofc do that first.

OpenStudy (anonymous):

ok so i got x-1, x+3, and x + 2

OpenStudy (anonymous):

i synthetically divided the inverse of each number and they all gave me a reminder of 0

Nnesha (nnesha):

looks good to me.

OpenStudy (anonymous):

Thank you so much for the help, you both were wonderful help! :) XD

Nnesha (nnesha):

np. good work btw keep it up!

OpenStudy (anonymous):

Ill go ahead and close the question now.